OpenStudy (anonymous):

Im using the logistic equation to work out a rabbit population question and just want to confirm my working out... dy/dt = ay(b - y) Separate variables: dy / ay(b-y) = dt Separate fractions: dy [1/aby + 1/ab(b-y)] = dt Multiply both sides by ab: dy [1/y + 1/(b - y)] = ab dt Integrate: ln(y) - ln(b-y) = abt + C Logarithm property: ln(y / b - y) = abt + C Algebra to solve for y: y / (b - y) = Ce^(abt) [positive C] -1 + b/b-y = Ce^(abt) b/b-y = Ce^(abt)+ 1 b - y = b / (Ce^(abt) + 1) y = b - b/(Ce^(abt) + 1) is this right?

OpenStudy (turingtest):

How did you separate the fractions? that part looks fishy

OpenStudy (turingtest):

...and if a is a constant I would leave it on the other side

OpenStudy (turingtest):

\[{dy\over y(b-y)}=adt\]then do partial fractions

OpenStudy (anonymous):

A rabbit population satisfies the logic equation \[dy/dt=2\times10^{-7}y(10^6-y)\] where t is the time measured in months. The population is suddenly reduced to 40% of its steady state size by myxamatiosis. If the myxamatosis then has no further effect, how large is the population 8 months later? How long will it take for the population to build up again to 90% of its steady state size?

OpenStudy (turingtest):

ok you just did it differently than me it looks good as far as I can tell

OpenStudy (anonymous):

so \[a=2\times10^-7\]and \[b=10^6\] so using the formula...\[dy/dt=ay(b-y)\] I need to take dt over to the RHS of the equation and bring everything else to the left?

OpenStudy (turingtest):

yeah but it seems to work out the same...\[{dy\over y(b-y)}\implies\frac1b(\frac1y-\frac1{b-y})=adt\implies \frac1y-\frac1{b-y}=abdt\]\[\ln {y\over(1-y)}=abt+C\]\[{y\over(b-y)}=Ce^{abt}=-1+{b\over b-y}=Ce^{abt}\]I just kept going for a morning exercise, but clearly you and I are going to get the same answer :)

OpenStudy (turingtest):

\[{b\over b-y}=Ce^{abt}+1\implies y=b-{b\over Ce^{abt}+1}\]yep :)

OpenStudy (anonymous):

awesome... thanks!!

OpenStudy (turingtest):

very welcome, thanks for the wake-up :)